# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ V.2 Stochastic gravitational wave background
$ds^{{[\rm ac]}2}\equiv-dt^{2}+\eta^{2}(t)dx^{i}dx_{i}=-\eta^{2}(\tau)\bigl{(}d\tau^{2}-g^{{[\rm ac]}}_{ij}dx^{i}dx^{j}\bigr{)}\,,$ (82)
@@
$\Box\bar{h}^{[\rm ac]}_{ij}(\mathbf{x},\tau)-2\frac{\eta^{{}^{\prime}}(\tau)}{\eta(\tau)}\bar{h}^{[\rm ac]^{\prime}}_{ij}(\mathbf{x},\tau)=16\pi GT^{[\rm ac]}_{ij}\,,$ (83)
@@
Following standard procedures, we introduce a Fourier transformation of this expression and we define $\tilde{h}^{[\rm ac]}_{\lambda}~{}\equiv~{}\eta(\tau)~{}h_{\lambda}$, so the field equation (83) may be recast in the form
${h}^{\prime\prime}_{\lambda}(\mathbf{k},\tau)+\Bigl{(}k^{2}-\frac{\eta^{\prime\prime}(\tau)}{\eta(\tau)}\Bigr{)}{h}_{\lambda}(\mathbf{k},\tau)=16\pi G\eta(\tau)T_{\lambda}(\mathbf{k},\tau)\,,$ (84)
@@
V.2.1 Sub-horizon condition with no source
Figure 12: Typical 3D solutions and contour plots of ${h}(\lambda,\mathbf{k},\tau)$ as a function of $\eta(\tau)$ and $\tau$ for a fixed values of $k$ and the dimensional constants. The results are presented in a pair of figures on the left, in accordance with equation (87), and in another pair of figures on the right, in line with equation (89).
@@ V.2.1 Sub-horizon condition with no source
$T^{[\rm ac]}_{\lambda}(\mathbf{k},\tau)\quad\rightarrow\quad p_{\lambda}(\eta(\tau))=-\eta(\tau)\eta^{\prime}(\tau)\,,$ (85)
@@
$\rightarrow{h^{\prime\prime}}_{\lambda}(\mathbf{k},\tau)+k^{2}{h}_{\lambda}(\mathbf{k},\tau)=-16\pi\eta^{3}\frac{{\eta^{\prime}}}{\eta}\,.$ (86)
In the standard formulation, when considering a generic scale factor obeying a power law such as $\eta(\tau)=\eta_{n}\tau^{n}$ [73], which covers the cases of radiation domination ($n=1$), matter domination ($n=2$), as well as De Sitter inflation ($n=-1$), the general solution of equation (83), with the source term set to zero, is given by [58, 73]
${h}(\lambda,\mathbf{k},\tau)=\frac{A_{\lambda}(\mathbf{k})}{\eta(\tau)}e^{ik\tau}+\frac{B_{\lambda}(\mathbf{k})}{\eta(\tau)}e^{-ik\tau},\quad\mbox{for}\quad k>>{\cal H}\,,$ (87)
@@
Figure 12 shows typical 3D solutions and contour plots of ${h}(\lambda,\mathbf{k},\tau)$ as a function of $\eta(\tau)$ and $\tau$ for fixed values of $k$ and the dimensional constants, based on equations (87) for the left pair of figures and equation (89) for the right pair of figures. The results presented in these figures depict systematic variations in the spacetime configuration of primordial matter due to the non-commutative formulation. These variations suggest a unique dynamics in the production of relic gravitational waves. Additionally, the color palette associated with the distribution of matter in these figures indicates greater intensity in the region prior to the expansion phase, which has predictable consequences for the acceleration of the Universe and the production of gravitational waves. It is worth noting that the solutions presented in the right figures are expected to remain constant in time for modes outside the Hubble radius [73]. However, these expectations may not align with the format of the present equation as well as with the evolutionary behavior of the dynamical equations.
V.2.2 Super-horizon condition with no source
Figure 13: Real and imaginary parts of 3D solutions and contour plots of the product $\eta(\tau){h}(\lambda,\mathbf{k},\tau)$ as a function of $\tau$ for different values of $n$, according to equation (90).
Figure 14: Real and imaginary parts of 3D solutions and contour plots of equation (90).
@@ V.2.2 Super-horizon condition with no source
$\rightarrow{h^{\prime\prime}}_{\lambda}(\mathbf{k},\tau)-\frac{\eta^{\prime\prime}(\tau)}{\eta}\tilde{h}_{\lambda}(\mathbf{k},\tau)=-16\pi\eta^{3}\frac{{\eta^{\prime}}}{\eta}\,.$ (88)
Assuming, as in the standard formulation, the power-law time dependence $\eta(\tau)=\eta_{n}\tau^{n}$ [73] and setting the source to zero, the typical solutions of equation (83) are as follows:
${h}(\lambda,\mathbf{k},\tau)=A_{\lambda}(\mathbf{k})+B_{\lambda}(\mathbf{k})\int_{0}^{\tau}\frac{d\tau}{\eta^{2}(\tau)}\quad\mbox{for}\quad k<<{\cal H}\,.$ (89)
V.2.3 General solution with no source
Indeed, due to the systematic contribution at the core of the differential equation from various terms related to the potential $V(\eta,\tau)$, we can make a good approximation by setting the source term to zero. This allows us to express the general solution of equation (83) as
${h}(\lambda,\mathbf{k},\tau)=\frac{A_{\lambda}(\mathbf{k})}{\eta(\tau)}\tau J_{n-1}(k\tau)+\frac{B_{\lambda}(\mathbf{k})}{\eta(\tau)}\tau Y_{n-1}(k\tau)\,,$ (90)
@@ V.2.3 General solution with no source
Figures 13 and 14 display typical sample solutions of equation (90). These figures provide insight into the behavior of $\eta(\tau){h}(\lambda,\mathbf{k},\tau)$ and ${h}(\lambda,\mathbf{k},\tau)$ under arbitrary parameter choices, focusing on their structural behavior. In figure 13, we observe the behavior of $\eta(\tau){h}(\lambda,\mathbf{k},\tau)$ as a function of $\tau$ for different values of $n$. In figure 14, we examine the behavior of ${h}(\lambda,\mathbf{k},\tau)$ as a function of $\eta(\tau)$ for a fixed value of $n$. These results suggest a turbulent and violent behavior of relic gravitational waves. It is important to note that the stochastic gravitational wave background, characterized by a homogeneous distribution, refers not to their primordial distribution but to their collective behavior at present. According to the authors of ref. [73], the integral in equation (89) decays as the Universe expands, indicating that gravitational waves are “frozen” outside the Hubble horizon, a mechanism similar to what occurs during inflation. Once they re-enter the horizon, tensor perturbations become sub-horizon modes again.
V.3 General solution in the presence of a source: super-horizon condition
For practical reasons during the calculations performed below, we have organized the following set of expressions based on equations (62), (63), (64), and (65):
$\frac{{\eta}^{\prime\prime}}{\eta}$ $=$ $-\frac{1}{2}\Bigl{(}\frac{\eta^{\prime}}{\eta}\Bigr{)}^{2}-\frac{3}{2}\frac{\alpha\gamma{\eta^{\prime}}}{\eta^{3\alpha+2}}-\frac{1}{2\eta^{2}}V(\eta,\tau)\,;$ (91)
@@ V.3 General solution in the presence of a source: super-horizon condition
$V(\eta,\tau)\!=\!g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}-\frac{3\alpha^{2}\tau}{\eta^{6\alpha-1}}\,.$ (92)
@@
$\frac{\eta^{\prime}}{\eta}=-\frac{1}{\eta}v(\eta,\tau)\,,$ (93)
@@
$\frac{1}{\eta}v(\eta,\tau)=\frac{1}{3\eta^{2}}\Biggl{\{}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{(}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{)}\tau-\frac{3}{2}\frac{\alpha^{2}\tau^{2}}{\eta^{6\alpha-1}}\Biggr{\}}\,.$ (94)
These equations describe the temporal evolution of the scale factor, $\eta(t)$, within the non-commutative branch-cut formulation, accounting for a first-order dependence on the time-derivative of $\eta(t)$. By combining equations (91) through (94), we can express equation (88) as
${h}^{{}^{\prime\prime}}(\tau)$$\Biggl{\{}\frac{1}{2}\frac{1}{9\eta^{4}}\Biggl{(}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{[}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{]}\tau-\frac{3}{2}\frac{\alpha^{2}\tau^{2}}{\eta^{6\alpha-1}}\Biggr{)}^{2}$ (95)
@@
The curves in figure 15 display partial solutions of equation (95) with the assumption $\eta(\tau)=\eta_{n}\tau^{n}$, with $n=1$. Starting from the left, we first consider only the presence of the quadratic term, and then systematically add each of the other terms individually, which are contained on the left side of the equation. In the final figure, we include both the quadratic term and the source term from the right side of the equation. The results suggest that, for $n=1$, it is a reasonable approximation to retain the dominant quadratic term and the source term while disregarding the other.
${h}^{{}^{\prime\prime}}(\tau)$ $+$ $\frac{1}{2}\frac{1}{9\eta^{4}}\Biggl{(}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{[}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{]}\tau-\frac{3}{2}\frac{\alpha^{2}t^{2}}{\eta^{6\alpha-1}}\Biggr{)}^{2}{h}(\tau)$ (96)
@@
This approach, in view of the formal difficulties in solving equation (95), allowed us to explore the parameter space adopted in the formulation more comprehensively. The pair of left figures in 16 depicts solutions of equation (96) corresponding to $n=2$, while the first figure on the right illustrates solutions of equation (96) considering the first two dominant terms corresponding to $n=-1$ The last figure represents a sample solution of equation (96). The results in this figure correspond to a homogeneous distribution of primordial gravitational waves. The sampling solutions consider the following boundary conditions: $\Psi(1)=1$, $\Psi^{\prime}(1)=0$ $\Psi(1)=0$, $\Psi^{\prime}(1)=1$.
Figure 15: The curves show partial solutions of equation (95) assuming $\eta(\tau)=\eta_{n}\tau^{n}$, with $n=1$. The figures illustrate the inclusion of terms in a stepwise manner, starting with the quadratic term and systematically adding each of the following terms separately from the left side of the equation. Finally, the last figure demonstrates the quadratic term together with the source term from the right side of the equation. The parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\alpha=1/2$; $\gamma=1$; ${\cal C}=1$; $n=1$.
Figure 16: The pair of figures on the left displays solutions of equation (96) corresponding to $n=2$. The first figure on the right shows solutions for $n=-1$. The last figure presents a sample solution of equation (95). The values of the parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\alpha=1/2$; $\gamma=1$; ${\cal C}=1$; $n=-1$.
V.4 Power spectrum and density parameter
@@
$\rho_{GW}=\langle T_{00}\rangle=\frac{c^{2}}{16\pi G}\Bigl{\langle}\sum_{\alpha}\Bigl{(}\frac{\partial h_{\alpha}}{\partial t}\Bigr{)}^{2}\Bigr{\rangle}\,,$ (98)
@@
Replacing (101) in (98) and comparing with (97), we obtain
$f\frac{d\rho_{GW}}{df}=\frac{\pi c^{2}}{4GT}f^{3}\sum_{\alpha}\Bigl{|}\tilde{h}_{\alpha}(t)\Bigr{|}^{2}\,.$ (102)
@@
$\Omega_{GW}=\frac{1}{\rho_{c}}\frac{d\,\rho_{GW}}{d\,\log f}=\frac{2\pi^{2}}{3H_{0}^{2}T}f^{3}\sum_{\alpha}\Bigl{|}\tilde{h}_{\alpha}(t)\Bigr{|}^{2}\,,$ (103)
@@
$\rho_{c0}=\frac{3c^{2}H_{0}^{2}}{8\pi G}\approx 7.6\times 10^{-9}erg/cm^{3}\,,$ (104)
@@
$S_{n}(f)=\frac{1}{2}f\sum_{\alpha}\Bigl{|}\tilde{h}_{\alpha}(f)\Bigr{|}^{2}\,,$ (105)
@@
$\rho_{GW}(k,\tau)$ $=$ $\int_{0}^{\infty}\frac{dk}{k}\frac{d\,\rho_{GW}}{d\,\log k}$ (108)
@@
$\frac{d\rho_{GW}}{d\,\log k}=\frac{k^{2}h^{2}(k,\tau)}{16\pi\eta^{2}(\tau)}=\Omega_{GW}(k,\tau)\,\rho_{c}\,.$ (109)
@@
$\Omega_{GW}(k,\tau)=\frac{1}{\rho_{c}}\frac{k^{2}h^{2}(k,\tau)}{16\pi\eta^{2}(\tau)}\,,$ (110)
@@
with $\rho_{c}$ defining the density closure of the Universe at time $\tau$ and $H^{2}=\Bigl{(}\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\Bigr{)}^{2}.$ From expressions (106) and (110) we obtain, for any time $\tau$:
$S_{n}(k,\tau)$ $=$ $\frac{3H^{2}(2\pi)}{k^{3}}\Omega_{GW}(k,\tau)=\frac{3H^{2}(2\pi)}{k^{3}}\frac{1}{\rho_{c}}\frac{k^{2}h^{2}(k,\tau)}{16\pi\eta^{2}(\tau)}$ (112)
@@
$\Omega_{GW}(f,\tau)=\frac{2\pi^{2}}{3H^{2}}f^{2}\frac{h^{2}(f,\tau)}{\eta^{2}(\tau)}\,,$ (113)
@@
$S_{n}(f,\tau)=\frac{1}{2f}\frac{h^{2}(f,\tau)}{\eta^{2}(\tau)}\,.$ (114)
VI Stochastic gravitational wave background

### 03 Feb 2024 v2

@@ 5.2 Stochastic gravitational wave background
$ds^{{[\rm ac]}2}\equiv-dt^{2}+\eta^{2}(t)dx^{i}dx_{i}=-\eta^{2}(\tau)\bigl{(}d\tau^{2}-g^{{[\rm ac]}}_{ij}dx^{i}dx^{j}\bigr{)}\,,$ (5.9)
@@
$\Box\bar{h}^{[\rm ac]}_{ij}(\mathbf{x},\tau)-2\frac{\eta^{{}^{\prime}}(\tau)}{\eta(\tau)}\bar{h}^{[\rm ac]^{\prime}}_{ij}(\mathbf{x},\tau)=16\pi GT^{[\rm ac]}_{ij}\,,$ (5.10)
@@
Following standard procedures, we introduce a Fourier transformation of this expression and we define $\tilde{h}^{[\rm ac]}_{\lambda}~{}\equiv~{}\eta(\tau)~{}h_{\lambda}$, so the field equation (5.10) may be recast in the form
${h}^{\prime\prime}_{\lambda}(\mathbf{k},\tau)+\Bigl{(}k^{2}-\frac{\eta^{\prime\prime}(\tau)}{\eta(\tau)}\Bigr{)}{h}_{\lambda}(\mathbf{k},\tau)=16\pi G\eta(\tau)T_{\lambda}(\mathbf{k},\tau)\,,$ (5.11)
@@
5.2.1 Sub-horizon condition with no source
Figure 12: Typical 3D solutions and contour plots of ${h}(\lambda,\mathbf{k},\tau)$ as a function of $\eta(\tau)$ and $\tau$ for a fixed values of $k$ and the dimensional constants. The results are presented in a pair of figures above, in accordance with equation (5.14), and in another pair of figures below, in line with equation (5.16).
@@ 5.2.1 Sub-horizon condition with no source
$T^{[\rm ac]}_{\lambda}(\mathbf{k},\tau)\quad\rightarrow\quad p_{\lambda}(\eta(\tau))=-\eta(\tau)\eta^{\prime}(\tau)\,,$ (5.12)
@@
$\rightarrow{h^{\prime\prime}}_{\lambda}(\mathbf{k},\tau)+k^{2}{h}_{\lambda}(\mathbf{k},\tau)=-16\pi\eta^{3}\frac{{\eta^{\prime}}}{\eta}\,.$ (5.13)
In the standard formulation, when considering a generic scale factor obeying a power law such as $\eta(\tau)=\eta_{n}\tau^{n}$ [72], which covers the cases of radiation domination ($n=1$), matter domination ($n=2$), as well as De Sitter inflation ($n=-1$), the general solution of equation (5.10), with the source term set to zero, is given by [58, 72]
${h}(\lambda,\mathbf{k},\tau)=\frac{A_{\lambda}(\mathbf{k})}{\eta(\tau)}e^{ik\tau}+\frac{B_{\lambda}(\mathbf{k})}{\eta(\tau)}e^{-ik\tau},\quad\mbox{for}\quad k>>{\cal H}\,,$ (5.14)
@@
Figure 12 shows typical 3D solutions and contour plots of ${h}(\lambda,\mathbf{k},\tau)$ as a function of $\eta(\tau)$ and $\tau$ for fixed values of $k$ and the dimensional constants, based on equations (5.14) for the left pair of figures and equation (5.16) for the right pair of figures. The results presented in these figures depict systematic variations in the spacetime configuration of primordial matter due to the non-commutative formulation. These variations suggest a unique dynamics in the production of relic gravitational waves. Additionally, the color palette associated with the distribution of matter in these figures indicates greater intensity in the region prior to the expansion phase, which has predictable consequences for the acceleration of the Universe and the production of gravitational waves. It is worth noting that the solutions presented in the right figures are expected to remain constant in time for modes outside the Hubble radius [72]. However, these expectations may not align with the format of the present equation as well as with the evolutionary behavior of the dynamical equations.
5.2.2 Super-horizon condition with no source
Figure 13: Real and imaginary parts of 3D solutions and contour plots of the product $\eta(\tau){h}(\lambda,\mathbf{k},\tau)$ as a function of $\tau$ for different values of $n$, according to equation (5.17).
Figure 14: Real and imaginary parts of 3D solutions and contour plots of equation (5.17).
@@ 5.2.2 Super-horizon condition with no source
$\rightarrow{h^{\prime\prime}}_{\lambda}(\mathbf{k},\tau)-\frac{\eta^{\prime\prime}(\tau)}{\eta}\tilde{h}_{\lambda}(\mathbf{k},\tau)=-16\pi\eta^{3}\frac{{\eta^{\prime}}}{\eta}\,.$ (5.15)
Assuming, as in the standard formulation, the power-law time dependence $\eta(\tau)=\eta_{n}\tau^{n}$ [72] and setting the source to zero, the typical solutions of equation (5.10) are as follows:
${h}(\lambda,\mathbf{k},\tau)=A_{\lambda}(\mathbf{k})+B_{\lambda}(\mathbf{k})\int_{0}^{\tau}\frac{d\tau}{\eta^{2}(\tau)}\quad\mbox{for}\quad k<<{\cal H}\,.$ (5.16)
5.2.3 General solution with no source
Indeed, due to the systematic contribution at the core of the differential equation from various terms related to the potential $V(\eta,\tau)$, we can make a good approximation by setting the source term to zero. This allows us to express the general solution of equation (5.10) as
${h}(\lambda,\mathbf{k},\tau)=\frac{A_{\lambda}(\mathbf{k})}{\eta(\tau)}\tau J_{n-1}(k\tau)+\frac{B_{\lambda}(\mathbf{k})}{\eta(\tau)}\tau Y_{n-1}(k\tau)\,,$ (5.17)
@@ 5.2.3 General solution with no source
Figures 13 and 14 display typical sample solutions of equation (5.17). These figures provide insight into the behavior of $\eta(\tau){h}(\lambda,\mathbf{k},\tau)$ and ${h}(\lambda,\mathbf{k},\tau)$ under arbitrary parameter choices, focusing on their structural behavior. In figure 13, we observe the behavior of $\eta(\tau){h}(\lambda,\mathbf{k},\tau)$ as a function of $\tau$ for different values of $n$. In figure 14, we examine the behavior of ${h}(\lambda,\mathbf{k},\tau)$ as a function of $\eta(\tau)$ for a fixed value of $n$. These results suggest a turbulent and violent behavior of relic gravitational waves. It is important to note that the stochastic gravitational wave background, characterized by a homogeneous distribution, refers not to their primordial distribution but to their collective behavior at present. According to the authors of ref. [72], the integral in equation (5.16) decays as the Universe expands, indicating that gravitational waves are “frozen” outside the Hubble horizon, a mechanism similar to what occurs during inflation. Once they re-enter the horizon, tensor perturbations become sub-horizon modes again.
5.3 General solution in the presence of a source: super-horizon condition
For practical reasons during the calculations performed below, we have organized the following set of expressions based on equations (4.23), (4.24), (4.25), and (4.26):
$\frac{{\eta}^{\prime\prime}}{\eta}$ $=$ $-\frac{1}{2}\Bigl{(}\frac{\eta^{\prime}}{\eta}\Bigr{)}^{2}-\frac{3}{2}\frac{\alpha\gamma{\eta^{\prime}}}{\eta^{3\alpha+2}}-\frac{1}{2\eta^{2}}V(\eta,\tau)\,;$ (5.18)
@@ 5.3 General solution in the presence of a source: super-horizon condition
$V(\eta,\tau)\!=\!g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}-\frac{3\alpha^{2}\tau}{\eta^{6\alpha-1}}\,.\!$ (5.19)
@@
$\frac{\eta^{\prime}}{\eta}=-\frac{1}{\eta}v(\eta,\tau)\,,$ (5.20)
@@
$\frac{1}{\eta}v(\eta,\tau)=\frac{1}{3\eta^{2}}\Biggl{\{}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{(}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{)}\tau-\frac{3}{2}\frac{\alpha^{2}\tau^{2}}{\eta^{6\alpha-1}}\Biggr{\}}\,.$ (5.21)
These equations describe the temporal evolution of the scale factor, $\eta(t)$, within the non-commutative branch-cut formulation, accounting for a first-order dependence on the time-derivative of $\eta(t)$. By combining equations (5.18) through (5.21), we can express equation (5.15) as
[…]
@@
The curves in figure 15 display partial solutions of equation (5.22) with the assumption $\eta(\tau)=\eta_{n}\tau^{n}$, with $n=1$. Starting from the left, we first consider only the presence of the quadratic term, and then systematically add each of the other terms individually, which are contained on the left side of the equation. In the final figure, we include both the quadratic term and the source term from the right side of the equation. The results suggest that, for $n=1$, it is a reasonable approximation to retain the dominant quadratic term and the source term while disregarding the other.
[…]
@@
This approach, in view of the formal difficulties in solving equation (5.22), allowed us to explore the parameter space adopted in the formulation more comprehensively. The pair of left figures in 16 depicts solutions of equation (LABEL:Mainmainmain) corresponding to $n=2$, while the first figure on the right illustrates solutions of equation (LABEL:Mainmainmain) considering the first two dominant terms corresponding to $n=-1$ The last figure represents a sample solution of equation (LABEL:Mainmainmain). The results in this figure correspond to a homogeneous distribution of primordial gravitational waves. The sampling solutions consider the following boundary conditions: $\Psi(1)=1$, $\Psi^{\prime}(1)=0$ $\Psi(1)=0$, $\Psi^{\prime}(1)=1$.
Figure 15: The curves show partial solutions of equation (5.22) assuming $\eta(\tau)=\eta_{n}\tau^{n}$, with $n=1$. The figures illustrate the inclusion of terms in a stepwise manner, starting with the quadratic term and systematically adding each of the following terms separately from the left side of the equation. Finally, the last figure demonstrates the quadratic term together with the source term from the right side of the equation. The parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\alpha=1/2$; $\gamma=1$; ${\cal C}=1$; $n=1$.
Figure 16: The pair of figures on the left displays solutions of equation (LABEL:Mainmainmain) corresponding to $n=2$. The first figure on the right shows solutions for $n=-1$. The last figure presents a sample solution of equation (5.22). The values of the parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\alpha=1/2$; $\gamma=1$; ${\cal C}=1$; $n=-1$.
5.4 Power spectrum and density parameter
@@
$\rho_{GW}=\langle T_{00}\rangle=\frac{c^{2}}{16\pi G}\Bigl{\langle}\sum_{\alpha}\Bigl{(}\frac{\partial h_{\alpha}}{\partial t}\Bigr{)}^{2}\Bigr{\rangle}\,,$ (5.25)
@@
Replacing (5.28) in (5.25) and comparing with (5.24), we obtain
$f\frac{d\rho_{GW}}{df}=\frac{\pi c^{2}}{4GT}f^{3}\sum_{\alpha}\Bigl{|}\tilde{h}_{\alpha}(t)\Bigr{|}^{2}\,.$ (5.29)
@@
$\Omega_{GW}=\frac{1}{\rho_{c}}\frac{d\,\rho_{GW}}{d\,\log f}=\frac{2\pi^{2}}{3H_{0}^{2}T}f^{3}\sum_{\alpha}\Bigl{|}\tilde{h}_{\alpha}(t)\Bigr{|}^{2}\,,$ (5.30)
@@
$\rho_{c0}=\frac{3c^{2}H_{0}^{2}}{8\pi G}\approx 7.6\times 10^{-9}erg/cm^{3}\,,$ (5.31)
@@
$S_{n}(f)=\frac{1}{2}f\sum_{\alpha}\Bigl{|}\tilde{h}_{\alpha}(f)\Bigr{|}^{2}\,,$ (5.32)
@@
$\rho_{GW}(k,\tau)$ $=$ $\int_{0}^{\infty}\frac{dk}{k}\frac{d\,\rho_{GW}}{d\,\log k}$ (5.35)
@@
$\frac{d\rho_{GW}}{d\,\log k}=\frac{k^{2}h^{2}(k,\tau)}{16\pi\eta^{2}(\tau)}=\Omega_{GW}(k,\tau)\,\rho_{c}\,.$ (5.36)
@@
$\Omega_{GW}(k,\tau)=\frac{1}{\rho_{c}}\frac{k^{2}h^{2}(k,\tau)}{16\pi\eta^{2}(\tau)}\,,$ (5.37)
@@
with $\rho_{c}$ defining the density closure of the Universe at time $\tau$ and $H^{2}=\Bigl{(}\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\Bigr{)}^{2}.$ From expressions (5.33) and (5.37) we obtain, for any time $\tau$:
$S_{n}(k,\tau)$ $=$ $\frac{3H^{2}(2\pi)}{k^{3}}\Omega_{GW}(k,\tau)=\frac{3H^{2}(2\pi)}{k^{3}}\frac{1}{\rho_{c}}\frac{k^{2}h^{2}(k,\tau)}{16\pi\eta^{2}(\tau)}$ (5.39)
@@
$\Omega_{GW}(f,\tau)=\frac{2\pi^{2}}{3H^{2}}f^{2}\frac{h^{2}(f,\tau)}{\eta^{2}(\tau)}\,,$ (5.40)
@@
$S_{n}(f,\tau)=\frac{1}{2f}\frac{h^{2}(f,\tau)}{\eta^{2}(\tau)}\,.$ (5.41)
6 Stochastic gravitational wave background